A RACE CAR HAS A MASS OF 710 KG. IT STARTS FROM REST AND TRAVELS 40 M IN 3 S. THE CAR IS UNIFORMLY ACCELERATED DURING THE ENTIRE TIME. WHAT NET FORCE IS APPLIED TO IT ?


1
Expert's answer
2022-02-17T04:34:07-0500

Given:

"m=710\:\rm kg"

"v_0=0"

"d=40\:\rm m"

"t=3\:\rm s"


The distance traveled by a car

"d=v_0t+at^2\/2=at^2\/2"

Hence, the acceleration of the car

"a=\frac{2d}{t^2}=\frac{2*40}{3^2}=8.89\:\rm m/s^2"

The Newton's second law says

"F=ma=710*8.89=6311\:\rm N"